# ConCodeEval: Evaluating Large Language Models for Code Constraints in Domain-Specific Languages

### v2

@@ Abstract
Recent work shows Large Language Models (LLMs) struggle to understand natural language constraints for various text generation tasks in zero- and few-shot settings. While, in the code domain, there is wide usage of constraints in code format to maintain the integrity of code written in Domain-Specific Languages [^1] https://w.wiki/6jCH (DSLs) like JSON and YAML which are widely used for system-level programming tasks in enterprises. Given that LLMs are increasingly used for system-level code tasks, evaluating if they can comprehend these code constraints is crucial. However, no work has been done to evaluate their controllability over code constraints. Hence, we introduce ConCodeEval first-of-its-kind benchmark having two novel tasks for code constraints across five representations. Our findings suggest that language models struggle with code constraints. Code languages that perform excellently for normal code tasks do not perform well when the same languages represent fine-grained constraints.
@@ 1 Introduction
Large language models (LLMs) have shown promising results (Brown et al., 2020) in generating coherent text and code in zero and few-shot settings, especially for resource-rich languages. However, their practical utility depends on their ability to follow instructions or constraints at various granularity, encompassing user and system requirements. Recent studies (Sun et al., 2023) indicate that LLMs struggle with fine-grained constraints in natural language prompts, especially in paraphrase generation and numerical planning tasks.
While previous work studied the controllability of language models (LMs) using fine-grained instructions in natural language (NL) format, the necessity to represent instructions in code format arises from a critical enterprise use case involving system-level programming. In enterprises, Listing 1: The JSON sample generated (highlighted in yellow) by the Granite 20B model does not adhere to the minContains and subsequent numerical constraints. Write a JSON sample with field values as per the JSON format schema given below. { "type": "array", "contains": { "type": "number", "multipleOf": 2.66, "exclusiveMinimum": 0.08231885995435284, "exclusiveMaximum": 5.1100233535478 }, "minContains": 7 } JSON sample: ‘‘‘ [2.66, 5.22, 8.88, 12.54, 16.2, 19.86, 23.52, 27.18] ‘‘‘ system-level code integrity is maintained through use-case-specific constraints, typically encoded using schemas. Schemas are instructions in structured code languages like JSON, YAML, XML, or Python to enforce constraints like data types and required fields, ensuring code integrity. For example, the schema in Listing LABEL:lst:task1 requires the sample to be an array of numbers. Each number must be a multiple of 2.66 and fall within the range defined by the exclusiveMinimum and exclusiveMaximum fields. Additionally, the array must contain at least seven elements. Following such schema constraints, developers write system-level code in Domain-Specific Languages (DSLs) in a format similar to JSON, YAML, or XML. DSLs are custom languages with specialized schemas and syntax suitable for a particular domain or application. These DSLs are common for tasks like data exchange and system configuration, such as in Kubernetes [^2] https://w.wiki/3kbZ. Writing DSL code requires deep domain expertise and a significant learning process for developers. This has led to a growing adoption of LLMs for system-level programming in several products such as Ansible Lightspeed [^3] https://developers.redhat.com/products/ansible/lightspeed.
Given the cruciality of factoring in schemas with LLMs, there is increasing interest in using constrained decoding for DSLs (Pimparkhede et al., 2024; Wang et al., 2024a). However, given its limitations (Appendix A.4), it is necessary to evaluate if LMs are cognizant of code constraints when directly presented as a part of the prompt. Therefore, we aim to study the controllability of LMs through two novel seed tasks: (i) Data as Code generation: valid sample generation factoring in constraints (ii) DSL validation: validate code against constraints. We evaluate two model families, Llama and Granite, ranging from 8B to 70B parameters, aligning with enterprise needs for system-level tasks, where open-source models provide an economical and transparent alternative to black-box models like GPT-4. Both tasks are highly motivated from research and enterprise use case point of view as detailed in Appendix A.5.
@@
    We introduce two novel NLP tasks for enterprise system-level code: code generation from fine-grained schema instructions and code validation against schemas. To the best of our knowledge, we are the first to evaluate LMs on these tasks.
@@
    Comparative and qualitative analysis of state-of-the-art code models for two novel tasks with different schema languages. Our findings show that language models best comprehend JSON schema (Tables 1, 2) and are agnostic of language proportions in pre-training data.
2 Data as Code Generation in DSL
Figure 1: Uniform trend of steep decline in performance across models for constraints positioned in the middle and beginning of the JSON schema context and output for task 1. We divide the schema in to $3$ equal portions Begin, Middle, and End, and put the violated constraints based on their locality into either of these three buckets.
Task description:
Given the schema, the generation task (see Listing LABEL:lst:task1) aims to produce a compliant data sample in DSL code format. We draw inspiration from several use cases (see Appendix A.5), including synthesizing schema-compliant data from LLMs’ parametric memory to train and evaluate smaller-sized models Song et al. (2020) and generating diverse sets of samples to be used in product test pipelines. For reliable DSL code generation, LLMs need to be schema-aware.
Dataset:
We synthetically prepare $602$ schemas for each of the $5$ representations having combinations of various constraints (Appendix A.6). First, we prepare JSON schemas using our combinatorial tool to generate a good mix of constraints. We then convert each JSON schema to XML and YAML schemas using automated tools. Further, we include resource-rich general-purpose language - Python using the Pydantic library generated using the Gemini-1.0-pro (Team et al., 2024) model as a code translation task. We extend our evaluation to NL representation generated using rule-based templates. We [^4] The schemas are manually validated by the paper’s authors. ensure equivalence of the generated schemas across languages. More details are in Appendix A.8.
Evaluation metric:
Experimental setup:
We report greedy decoding results since it performed slightly better than beam search with a beam width of 3. More details including hyper-parameters in Appendix A.7.
Prompts:
We experiment with zero-shot and 3-shot prompting for each model. For 3-shot prompting, we identify errors from the zero-shot setting, then select shots similar to the most frequent errors. Examples of prompts are in Appendix LABEL:lst:task1. While we represent the zero-shot results in the main paper, few-shot results are in the Appendix (Table 2).
Results:
Among the $5$ schema representations studied, NL is best understood by models across all outputs. JSON and YAML schemas perform well for constraints in code despite their limited presence in pre-training data. Surprisingly, models struggle with constraints in Python, though it is the major portion of the pre-training data, and models also find XML schemas challenging. Notably, models did not exhibit a performance boost when schema and output representation languages were the same. As shown in Figure 1, models are sensitive to the locality of the constraints in the schema and struggle to factor in constraints present in the beginning and middle portions of the schema. Irrespective of their overall performance, models show a similar distribution of mistakes across all constraint types (see Table 4) and show improved performance in few-shot results (see Table 2) when shots relevant to the mistakes used. Among the two family of models, Llama3 70B performed the best followed by Granite 34B.
@@ Results:
3 DSL Validation
Task description:
There is a growing body of work Hada et al. (2024) on showing promising usage of LLMs as evaluators in many tasks. On similar lines, given the DSL sample and schema to validate, this task (see Listing LABEL:lst:task2) aims to determine the validity of the provided sample against the constraints through boolean question answering (QA). Also, the task is highly motivated from various use cases (see Appendix A.5) and throws light on LM’s understanding of the relation between requirements and output in various representations.
@@ 3 DSL Validation
For each of the $602$ schemas across $5$ representations as described in Section 2, we generate $3$ data samples across JSON, XML, and YAML languages. First, these data samples are synthetically generated by parsing through the JSON schema by randomly pruning and selecting constraints, resulting in data samples of different lengths and constraints. We then convert the generated JSON data samples to equivalent YAML and XML formats. Dataset consists of $3076$ instances with $45\%$ of no and $55\%$ yes instances.
@@
Since it is a boolean QA task, we use Macro average F1 (see Table 6) and Accuracy (Val Acc) as evaluation metrics (see Table 1).
@@
The decoding strategy used here is similar to the generation task as mentioned in Section 2. More details in Appendix A.7.
@@
The goal of this task is to answer with either yes or no. We experiment with zero- and few-shot prompting. With few shot prompting we provide one example each of yes and no answer. Results for few-shot prompting and examples of prompts are given in Appendix (Table 2).
Results:
@@
Although NL representation excels in generation tasks, it degrades the validation performance of larger models like 70B. JSON, YAML and Python representations show effectiveness in one of the output formats however poorly perform for others. Like in task 1, models perform sub-optimally when both schema and output representations are same. In lines with task 1, XML stands as a challenging language for models. The Llama3 70B model perform best in validation like in task 1, with other models hovering around 50% Val Acc, likely reflecting random choice given the binary nature of the task. Smaller models, particularly the Llama3-8B with natural language representation, show notable improvement, as its pre-training is a combination of NL and code.
@@ 4 Related Work
There is extensive work Muennighoff et al. (2024); Cassano et al. (2022) on evaluating capabilities of LLMs for various code tasks such as code completion, translation, etc for resource-rich languages like Python. Despite there being work Cassano et al. (2022) on multi-lingual code, there is scant attention to low-resource languages such as DSLs, though having crucial importance. In parallel, using LLMs as evaluators for low-resource languages is gaining interest, however limited, mainly focusing on languages like XML and INI Lian et al. (2024).
@@ 5 Conclusion
We introduce two novel tasks - Data as Code generation and DSL validation to test the controllability of LLMs when constraints are in code format, which is a crucial use case for system-level programming tasks in enterprises. We evaluate LLMs over $5$ schema representations, including YAML, JSON, Python, XML, and NL, and $3$ output representations, including YAML, JSON, and XML. We conclude that model performance does not directly correlate with language’s portion in pre-training data. Models for task 1 best understand NL. However, this is not the case with the validation task. Interestingly, models underperformed when the schema and output representations were the same, and the locality of the constraints in the schema impacted their performance. Task 2 results show that most of the models, irrespective of their size, found it very challenging since they performed just above or under the 50% accuracy. We hope our work can help innovation in improving the capabilities of LLMs for such challenging use cases and serve as a valuable reference.

### v3

@@ Abstract
System-level programming is essential for modern enterprise infrastructure, enabling the automation and management of complex systems through declarative code. Developers write this code based on schemas, which themselves are a form of code that defines constraints like data types and required fields. These schemas help ensure operational correctness and smooth integration across systems. However, as enterprise schemas become complex, manually writing code adhering to these constraints becomes challenging for developers. Large Language Models (LLMs) have demonstrated potential in code generation and natural language understanding, particularly in zero-shot and few-shot settings. However, applying LLMs to handle constraints represented in code, essential for system-level programming rather than natural language, has not been explored. Hence, we introduce ConCodeEval, a study across two key dimensions: format and constraint efficacy with a first-of-its-kind benchmark involving two novel experiments for code constraints across five representations (JSON, YAML, XML, Python, and natural language). Our findings suggest that conscious use of representations can lead to optimal use of LLMs in enterprise use cases involving constraints. Nonetheless, LLMs still struggle with code constraints motivating the need for innovation in this direction.
@@ 1 Introduction
System-level programming is the backbone of modern enterprise infrastructure, enabling developers to define, manage, and automate complex systems seamlessly. Numerous enterprises use concepts like Infrastructure as Code [^1] https://en.wikipedia.org/wiki/Infrastructure_as_code (IaC) to let developers write declarative code. Such code must adhere to constraints called schemas, which define rules, including data types, required fields, and valid value ranges, ensuring operational correctness and smooth integration. For instance, the schema in Listing LABEL:lst:task1 mandates an array of even numbers within specific bounds, containing 1 to 7 elements.
Listing 1: The JSON sample generated (highlighted in yellow) by the Granite 20B model does not adhere to the minContains and subsequent numerical constraints specified in the schema.
Write a JSON sample with field values as per the JSON format schema given below.
"type": "array",
"contains": {
"type": "number",
"multipleOf": 2,
"exclusiveMinimum": 0,
"exclusiveMaximum": 65535
"minContains": 1,
"maxContains": 7
JSON sample:
[2, 3, 4, 6, 8, 10, 12, 14]
Schemas are crucial in real-world enterprise settings. For instance, deploying a database service in an OpenShift cluster involves writing compliant code with the correct attributes, such as the number of instances, port number to expose, compute to allocate, etc. Developers write system-level code in structured Domain Specific Languages (DSLs) such as JSON, YAML, XML, or Python, adhering to strict schema constraints. However, enterprise schemas are often complex and difficult to learn, slowing development and increasing errors. As a result, the need for automated and accurate systems for system-level programming is increasing leading to products such as Ansible Lightspeed Lig .
LLMs have shown great promise in generating coherent text and code in zero-shot and few-shot settings, making them highly appealing for system-level coding (Brown et al., 2020; Roziere et al., 2023; Mishra et al., 2024). Using LLMs to handle constraints represented in natural language (NL) has been extensively explored for tasks like poem generation and summarization Sun et al. (2023). However, Unlike these natural language tasks, constraints are often represented as code for system-level programming; hence, evaluating LLMs requires a different approach. In addition to assessing how well models adhere to constraints expressed in natural language, we must examine their ability to process, interpret, and generate structured formats while ensuring schema compliance. To ensure this, we evaluate LLMs under two key dimensions: Format Efficacy and Constraint Efficacy.
Format efficacy involves studying the performance of LLMs on varying constraint representations that form the input and output representations downstream enterprise use cases can consume. Specifically, we aim to answer the following research questions (RQ) for format efficacy: 1) Which format is optimally suited for constraint and output representation? 2) What is the trade-off between performance and context length cost? While constraint efficacy involves studying LLMs’ performance on various schema constraints within a format. Precisely, we aim to answer the following research questions related to constraint efficacy: 1) How does performance vary across different types of constraints? 2) What are the ideal positions for constraints in the schema for better adherence?
We prepare first-of-its-kind benchmark test set and conduct two experiments involving $5$ schema formats (JSON, YAML, XML, Python, and NL) and $3$ output formats (JSON, YAML, and XML) resulting $15$ combinations of use cases to investigate the aforementioned research questions. 1) Data as Code Generation (Section 2.1. 2) Data Validation (Section 2.2). This study provides insights into leveraging LLMs effectively for system-level programming tasks involving code constraints in enterprises.
@@
    First-of-its-kind study of language models for crucial industry use case of system-level programming involving code format constraints across four key dimensions: Format and Constraint efficacy.
@@
    Comparative and qualitative analysis of state-of-the-art language models involving code generation from fine-grained schema instructions and code validation against schemas. To the best of our knowledge, we are the first to evaluate LLMs code constraint competency.
2 Experiments
2.1 Data as Code Generation in DSL
Figure 1: Uniform trend of steep decline in performance across models for constraints positioned in the middle and beginning of the JSON schema context and output for data as code generation experiment. We divide the schema into $3$ portions, Begin, Middle, and End, and put the violated constraints based on their locality into either of these three buckets.
@@ 2.1 Data as Code Generation in DSL
Description:
Given the schema, the experiment (see Listing LABEL:lst:task1) aims to produce a compliant data sample in DSL code format. We draw inspiration from several use cases (see Appendix A.3), including synthesizing schema-compliant data from LLMs’ parametric memory to train and evaluate smaller-sized models Song et al. (2020) and generating diverse sets of samples to be used in product test pipelines. For reliable DSL code generation, LLMs need to be schema-aware.
@@
We synthetically prepare $602$ schemas for each of the $5$ representations having combinations of various constraints (Appendix A.4). First, we prepare JSON schemas using our combinatorial tool to generate a good mix of constraints. A combinatorial data generation tool factors in constraints of interest, constraint-specific information, and combinatorial preferences to generate the schemas. We then convert each JSON schema to XML and YAML schemas using openly available automatic lossless language-to-language translation tools. Further, we include resource-rich general-purpose language - Python using the Pydantic library generated using the Gemini-1.0-pro (Team et al., 2023) model as a code translation task. We extend our evaluation to NL representation generated using rule-based templates. We [^2] The schemas are manually validated by the paper’s authors. ensure equivalence of the generated schemas across languages. We plan to open-source all the scripts used for data preparation. Table 5 gives details regarding schema token length.
Evaluation metric:
Experimental setup:
We report greedy decoding results since it performed slightly better than beam search with a beam width of 3. We perform inference for all the models in $bfloat16$ precision and a max new token limit of 1024 tokens.
Prompts:
We experiment with zero- and 3-shot prompting for each model. For 3-shot prompting, we identify errors from the zero-shot setting, then select shots similar to the most frequent errors. We observe that most errors made by all the models are regarding short schema and the schema having root type of array as shown in sample LABEL:lst:task1. An example of a 3-shot prompt for a DSL generation experiment is shown below. Examples of prompts are in Appendix LABEL:lst:task1.
Output Representation
JSON YAML XML
Model Schema Gen Acc Val Acc Gen Acc Val Acc Gen Acc Val Acc
Llama3 8B JSON 48.3 71.2 46.6 68.1 39.2 64.1
Granite 8B 51.2 69.2 52.3 66.1 47.8 65.8
Granite 20B 58.3 73.5 56.4 72.3 50.2 68.2
Granite 34B 66.3 76.2 64.5 75.4 51.3 73.2
Codellama 34B 65.1 75.1 63.4 73.2 50.6 71.2
[…]
Llama3 8B XML 46.6 65.8 42.3 63.4 36.6 60.1
Granite 8B 46.2 64.8 44.5 63.2 34.5 57.3
Granite 20B 50.4 66.7 48.2 64.1 36.4 56.1
Granite 34B 52.3 68.5 51.1 63.4 39.2 53.2
Codellama 34B 49.2 66.2 49.2 63.2 35.1 52.1
[…]
Llama3 8B YAML 46.7 67.2 45.3 64.2 43.5 63.2
Granite 8B 48.1 65.2 46.2 61.2 44.2 61.2
Granite 20B 52.3 68.9 49.7 66.7 47.8 65.1
Granite 34B 54.2 67.7 51.3 65.3 45.3 56.4
Codellama 34B 56.8 66.4 50.2 64.3 47.8 56.2
[…]
Llama3 8B Python 43.2 60.1 41.1 58.9 39.2 57.6
Granite 8B 45.1 60.5 46.7 59.4 37.4 56.0
Granite 20B 48.2 57.2 45.9 57.8 38.4 58.2
Granite 34B 50.6 59.2 47.1 55.6 41.3 57.3
Codellama 34B 47.2 56.4 45.3 57.2 39.2 55.1
[…]
2.2 DSL Validation
Description:
There is a growing body of work Hada et al. (2024) on showing promising usage of LLMs as evaluators in many tasks. On similar lines, given the DSL sample and schema to validate, this experiment (see Listing LABEL:lst:task2) aims to determine the validity of the provided sample against the constraints through boolean question answering (QA). Also, the experiment is highly motivated from various use cases (see Appendix A.3) and throws light on LM’s understanding of the relation between requirements and output in various representations.
Dataset:
We synthetically prepare $602$ schemas across $5$ representations having combinations of hard and soft constraints. First, we prepare JSON schemas using our combinatorial tool to generate a good mix of constraints. We then convert each JSON schema to XML and YAML schemas using automated tools to ensure equivalence across representations. Further, we include Python representation using the Pydantic library as a resource-rich general-purpose language in our evaluation generated using the Gemini-1.0-pro (Team et al., 2023) model as a code translation task. We extend our evaluation to natural language representation generated using rule-based templates over the JSON schema. We [^3] The generated Python samples are manually validated by the paper’s authors. ensure equivalence of the generated schemas across languages by manually eyeballing the samples.
@@ 2.2 DSL Validation
Since it is a boolean QA experiment, we use Macro average F1 (see Table 6) and Accuracy (Val Acc) as evaluation metrics (see Table 1).
@@
The decoding strategy used here is similar to the data generation experiment as mentioned in Section 2.1. We perform inference in $bfloat16$ precision and a max new token limit of 1024 tokens. For beam search decoding, we use the beam width of 3.
@@
The goal of this experiment is to answer yes or no. We experiment with zero- and few-shot prompting. With few shot prompting, we provide one example each of yes and no answers. Results for few-shot prompting and examples of prompts are given in Appendix (Table 2).
3 Format Efficacy
3.1 Objective
To identify the most effective schema representation and output format for system-level programming while employing language models. Since schemas can be represented in various structured formats, including JSON, YAML, XML, Python, and even NL, determining which format best enables constraint adherence for language models while balancing context-length costs is critical.
3.2 RQ1: Which format is optimally suited for constraint and output representation?
Finding 1.
In the data as code generation experiment (section 2.1), models best understand (Table 1) NL across all outputs. At the same time, JSON and YAML schemas perform well (Table 2) for constraints in code despite their limited presence in pre-training data. Surprisingly, models struggle with constraints in Python, likely due to a bias toward generating general-purpose Python code rather than schema-specific patterns. In contrast, JSON and YAML schemas benefit from their rigid structures and alignment with schema-centric applications, making them easier for models to interpret.
Finding 2.
Using the same schema and output representation does not always enhance performance. For instance, in Table 2, YAML as schema and JSON as output representation performed better than YAML for both representations.
Finding 3.
@@ 3.2 RQ1: Which format is optimally suited for constraint and output representation?
Although NL representation excels in generation experiments, it degrades the validation performance of larger models like 70B. Like generation experiment, models perform sub-optimally when schema and output representations are the same. In line with the first experiment, XML stands as a challenging language for models. The Llama3 70B model performs best in validation as in the first experiment, with other models hovering around 50% Val Acc, likely reflecting the random choice given the binary nature of the experiment. Smaller models, particularly the Llama3-8B with natural language representation, show notable improvement, as its pre-training combines NL and code.
Key takeaway.
NL is a favorable language for schema representation, however, since its possible that enterprises lean more toward structured languages for better interoperability in which case JSON and YAML are ideal candidates for schema representation with JSON being favourable candidate for output representation. Nonetheless, the inconsistency in performance across experiments and model sizes underscores need for better schema comprehension and improved training strategies for NLP tasks involving validation.
3.3 RQ2: What is the trade-off between performance and context length cost?
Findings.
From section 3.2 key takeaway, JSON and YAML are ideal candidates for schema representation which form the context to the LLM. From Table 5, representing schema in YAML on an average takes $\sim$35less tokens than JSON. However, while choosing YAML would mean taking a drop of $\sim$14in Gen Acc and $\sim$4in Val Acc performance compared to JSON.
Key takeaway.
Enterprises should be cognizant of such tradeoff and choose ideal representation that fits their use case. Further, better tokenizer training techniques might lead to lower token expenditure for the desired representation.
4 Constraint Efficacy
4.1 Objective
To examine how language models handle various types of constraints embedded within schemas. Enterprise schemas enforce structural (e.g., required fields, data types) and semantic (e.g., dependencies, value constraints) rules.
4.2 RQ1: How does performance vary across different types of constraints?
Findings.
The analysis of the results shows that LLaMA3 8B and 70B exhibit similar patterns of missing constraints when generating JSON samples from a given schema (Table 3). In particular, constraints such as type, multiple, and exclusiveMinimum are often missing, while constraints such as maximum, additionalProperties, and minimum are more frequently followed. The high error rate in fundamental constraints type can be because training data contains many JSON-like samples where type is implicit rather than explicitly stated. The reason behind missing constraints like exclusiveMinimum and multipleOf may be because they involve high numerical precision. LLMs treat numbers as tokens, leading to potential rounding errors or incorrect enforcement.
Key takeaway.
LLMs struggle with numerical constraints underscoring need for better techniques throughout the stack from tokenizer to training. For enterprises, a rudimentary solution is to integrate constrained decoding or use post-processing validation to correct missing constraints after generation.
4.3 RQ2: What are the ideal positions for constraints in the schema for better adherence?
Findings.
We categorize the constraints of the schema into three sections based on tokens: beginning (first 30%), middle (next 40%), and end (last 30%). Later, we perform a needle-in-the-haystack experiment for the data-as-code generation. The heatmap in Figure 1 shows the statistics of constraints missed at every position for JSON to JSON generation. It reveals a consistent trend where models struggle the most with constraints positioned at the beginning of the schema, followed by the middle. In contrast, constraints at the end are least frequently missed. This suggests that models may prioritize constraints appearing later in the schema, likely due to the left-to-right decoding nature of autoregressive models, causing early constraints to be overwritten or ignored. We also observe that constraints in the middle position of the schema are frequently missed. This aligns with previous findings that the middle part of the long context is often missed Liu et al. (2024). For the data validation task, we analyze attention maps, which reveal a similar trend where the model pays less attention to the middle part of the schema (Figure 2).
Key takeaway.
This suggests that important constraints should be placed at the end of the schema or the beginning for longer schemas, depending on the use case.
@@ 5 Related Work
There is extensive work Muennighoff et al. (2024); Cassano et al. (2023) on evaluating capabilities of LLMs for various code tasks such as code completion, translation, etc, for resource-rich languages like Python. Despite there being work Cassano et al. (2023) on multi-lingual code, there is scant attention to low-resource languages such as DSLs, though having crucial importance. One notable work He et al. (2024), studies the bearing of prompt format in DSLs with LLM performance, however, does not include impact of output formats and controllability aspect in terms of code constraints crucial for enterprises. Further, using LLMs as evaluators for low-resource languages is gaining interest, however limited, mainly focusing on languages like XML and INI Lian et al. (2023).
@@ 6 Conclusion
We evaluate LLMs for system-level programming across two key dimensions: Format Efficacy and Constraint Efficacy. Format efficacy examines how LLMs handle different constraint formats, while constraint efficacy assesses their performance on various schema constraints within a format. We conduct two novel experiments to study these aspects: Data as Code generation and DSL validation. We evaluate LLMs across $5$ schema(YAML, JSON, Python, XML, NL) and $3$ output formats(YAML, JSON, XML). Our findings reveal that model performance does not directly correlate with a language’s presence in pre-training data. JSON and YAML are best suited for system-level programming, and enterprises should convert Python and XML formats to one of these for better LLM performance. We also observe that schema constraint locality affects performance, with constraints in the start and middle being most frequently violated. Placing critical constraints at the end improves reliability. We hope our work drives innovation in improving LLM capabilities for crucial industry use case of system-level programming involving code constraints.